a teacher wants to earn $1,500 per year in supplemental income from a cash gift of $16,000. she puts $6,000 in a bank that pays 7% annual interest. What rate must she earn on the remainder to achive her goal
Thank you
Vanessa

Answer by edjones(8007)

let x=rate of interest on the remainder.
6000*.07+10000x=1500
420+1000x=1500
subtract 420 from each side. 10000x=1500-420
10000x=1080
divide 10000 into each side. x=1080/10000=.108=10.8%
Check:
420+10000*.108==420+1080=1500
Ed